OpenStudy (anonymous):

80=10〖log〗_10*x/〖10〗^(-12) How do I find x? Do I simply times by x then divide by 80?

OpenStudy (anonymous):

It just means, 10log_10, I've been taught to type it out like that.

OpenStudy (anonymous):

so 〖log〗= log

OpenStudy (loser66):

ha, TJ here, you are safe. wait for him

terenzreignz (terenzreignz):

That's not fair, what if I only came to watch? :3 Oh well, anyway.... ^_^

terenzreignz (terenzreignz):

We need to remove all ambiguities, is THIS your equation? \[\Large80 = \frac{10\log_{10}x}{10^{-12}}\]

OpenStudy (anonymous):

10^-12 is over x and x is not apart of the log. Sorry for the misunderstanding and my poor explanation skills

terenzreignz (terenzreignz):

No worries :) Mathematics is a universal language, but somehow, translations get messy... We have to establish what your equation is first... \[\Large80 = \frac{10\log10}{\frac{10^{-12}}x`}\]

terenzreignz (terenzreignz):

This?

OpenStudy (anonymous):

Haha, sorry no no. it's 10log_10 * x/10^-12

terenzreignz (terenzreignz):

If all else fails, you could just use the draw option down there

terenzreignz (terenzreignz):

oh, I see... \[\Large \Large80 = \frac{[10\log10]\cdot x}{10^{-12}}\]

terenzreignz (terenzreignz):

That it?

OpenStudy (anonymous):